OpenStudy (anonymous):

f(x)=e^(x+4lnx) find f'(6).

OpenStudy (cathyangs):

that's about 871,406.2

OpenStudy (anonymous):

can your work and i need 2 decimal places

OpenStudy (anonymous):

*show

OpenStudy (cathyangs):

*sigh.* I'm just a mite bit busy with my hw...but the decimals are .19, I'm pretty sure.

OpenStudy (anonymous):

thats not the answer... ive been having trouble with problem all day.

OpenStudy (anonymous):

*this

OpenStudy (bahrom7893):

f(x)=e^(x+4lnx) f'(x) = e^(x+4lnx) * (x+4lnx)' (x+4lnx)' = 1 + 4/x f ' (x) = e^(x+4lnx) * (1 + 4/x)

OpenStudy (bahrom7893):

f'(6) = e^(6+4ln6) * (1 + 4/6) = e^(6+4ln6) * (10/6)

OpenStudy (bahrom7893):

f'(6) = e^(6+4*ln6) * (10/6) = 2160e^6

OpenStudy (anonymous):

o my god! that is correct!

OpenStudy (anonymous):

THANKS!

OpenStudy (bahrom7893):

np

OpenStudy (turingtest):

dude, really? typo the whole time?

OpenStudy (anonymous):

lol not really. i never use parenthesis.

OpenStudy (bahrom7893):

i was copy pasting turing.. where's the typo?

OpenStudy (anonymous):

he was talking to me

OpenStudy (bahrom7893):

ohh ok